The lowest form of the product  2 whole 3/7 × (7/9) is                                                          

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The lowest form of the product 2 whole 3/7 × (7/9) is (17/9) or 1 whole 8/9

First we have to convert the mixed fraction into improper fraction 2 whole 3/7 = 17/7 Then, (17/7) × (7/9)

= 17/9
= 1 whole 8/9


Thanks